A molding machine prepares a certain kind of car spare part with a target



diameter p = 42.23 mm. The machine has some variability, so the standard



deviation of the diameter is o = 0.003 mm. A sample of 10 spare parts is



inspected each hour for the process control purposes and record are kept of



the sample mean diameter. What will be the mean and standard deviation of



the number recorded?

1
Expert's answer
2022-03-30T04:25:46-0400

"\mu _{\bar{x}}=\mu =42.23\\\sigma _{\bar{x}}=\frac{\sigma}{\sqrt{n}}=\frac{0.003}{\sqrt{10}}=0.000948683"